Archive for May 2016

Apiary Report: 29th May. Health Check by SBI – the dreaded EFB confirmed

All was going so well , starting from the furthest colonies from the entrance, Top bar, Warre, No# 8, ect until we got to No3 .  All went pear shaped.     Debbie Burton ,the Seasonal Bee Inspector, discovered a few cells of EFB.  This colony had been small from the start of the season and slow to build up,  Debbie thought that the EFB may have been present over winter and not shown itself until now.    For those present it was a very useful experience to see the signs of EFB at first hand and what to look for in the brood frames.    If proof were needed, do not try to keep a failing colony going and certainly do not combine it with another colony unless you are sure that it has no disease.   Once condemned ,the colony had to be culled ASAP later that evening, when as many flying bees had returned as possible and the frames and bees burnt later.    Debbie advised that the boxes, floor, Qx, crown board  and roof would need to be scorched with a burner and removed to the container, for possible use next year.    A very sad end to the day, but necessary for the health of the other colonies.

Temperature  20 degrees and sunny.  Little wind.  All hives inspected by Debbie Burton  SBI  working from 15H back to No1

Hive#1   Brood chamber:     14 x 12,   QNS,  no Qc’s,   5 f  BAS,   stores ok, room ok. 1 super.   temper 10.   It was hoped to find the Queen and kill, as the colony is suffering very badly with chalk brood. – Next time…

Hive#2    Brood Chamber:    14 x 12,    QNS,  no Qc’s,  8f BAS,     stores ok, room ok,  1 super,  temper 8.  Wax moth grub found and killed, some bald brood evident, but Q laying well.

Hive#3    Brood Chamber: National Deep.  QNS, 3 f  BAS, – Ants present in hive, bees “jittery”, poss chronic bee paralysis,  then grey misshapen larva discovered, tested for EFB and confirmed positive.  B****r.   To Cull bees after flying bees return by sealing hive up and dosing with petrol.   Frames and bees to be burnt later. Apiary now on lockdown,  Nothing to be removed from site.

Hive#4    Brood Chamber:  Deep and Half,   QS,  9 Qc’s,  12 f BAS,  stores ok, room ok, 2 supers.   Some Qc’s sealed but Q still present.  Frame of brood with 2 Qc’s and f- food put into poly Nuc,with some extra bees to make up a new colony.  Put on stand in fern area away from trees.  Remaining  Qc’s taken down, and removed frames replaced with foundation to give more room.  Super frame with lots of drone cells removed for varroa check, replaced with DC.

Hive#5   Brood Chamber:  Deep and Half.    QS,   2 Qc’s one sealed.  needed more room so moved end frames into brood nest,  good distribution of new brood in shallow , put stores  up into Super to give more room in Shallow.

Hive#6    Brood Chamber:   Deep and Half,   QNS.   no Qc’s,   8 f BAS,  stores ok, needs more room, day old eggs, but possible ready to swarm, moved shallow f, into brood nest to give more room.  Wax  moth present in f 5 and 8.

Hive#7    Brood Chamber:   New Deep.  QNS,  12 charged Qc’s,   No brood, stores ok, room ok.  Artificial swarm made on 18/5.  but Q may have gone.  2 Qc’s left, others taken down.  New super of DC added as Super full.

Hive#7 A  Brood Chamber:   4 Shallows,  QNS,  12 Qc’s     8 f BAS,  stores ok, room ok. 1 super added.   Ants coming in.  1 sealed Qc -small, opened and larva dead. all brood sealed. QC opened so may be virgin Q ,  other Qc’s taken down.

Hive#8    Brood Chamber:   National Deep.  QS,  7 Qc’s,   8 f BAS,   stores ok, needs room,  Qc’s charged,  drones being laid in worker cells, Chalk brood present, sbi advised Regicide (Q killed)  widespread pepperpot brood. sealed QC left on F6.

Hive#9    Warre Hive.    2 brood chambers,  QS,  no Qc’s, 2 f BAS,  not much stores, small colony, old comb needs replacing if we are to keep going with colony.  No sign of any disease,   few ants around.  Solid floor may have caused mouldy frames, poss due to lack of ventilation over winter.

Hive#10    Top Bar Hive.   QNS, no Qc’s,  3 f BAS, little stores, room ok.     Wax moth and bald brood present.   one cell found with 2 eggs in!  Building but still weak. – ( needs swarm dumped in , as per Peter Reader)

Hive#15F   Brood Chamber:  National Deep. QS,  no Qc’s,  4 f BAS,   Sac brood (early stages) present on f6. some bald brood also present, along with varroa damage.

Hive#15H   Brood Chamber:  National Deep,   QS,  no Qc’s,  4 f BAS,  varroa damage, Sac brood and Bald brood present.  Laying drones in worker cells. stores ok but short of room, building wild comb in roof above cb.  A super of DC was added to help.

 

 

 

Next Apiary Meeting: Sunday 29th May 14:00, Health check

Annual health check by the seasonal bee inspector, Debbie Burton.    She will be going through the hives pointing out any problems and showing us all what to look for in our hives to keep them free of disease.

Richard, Liz and Mel,     Apiary custodians.

Apiary Report: Sunday 15th May.

Another lovely Sunday with lots to do.  Now is the time that the colonies are getting going, and its the ideal time to get hands on experience on colony manipulations.    Lots of activity in the hives, and all bar one or two making up nicely.    Nectar and pollen incoming, thanks to the warm weather and the Queens laying up to nearly her maximum, ready for the main flows in early June.        Number one priority was to go through the Warre hive for the first time this year.   What a mess,  lots of unattatched comb,  and floor debris , but the bees were still there and surviving.  The problems of operating this type of hive were apparent and though mimicing the natural hollow tree, it is not easy to check through regularly and as for swarm controls, very difficult.

Temperature   16 degrees C,    Sunny periods.  little wind.

Hive#1    Inspected by Brian Herbert

Brood  Chamber:  14 x 12,    QS,  Play cups,  6f Brood at All Stages,  (BAS)  3f stores,  room ok.  1 super . empty. Temper 7,   vb out.  –   Lots of chalkbrood,  A shallow frame was put into broodnest for drone trapping part of IPM.       Recomended  to requeen ASAP  to overcome chalkbrood.

Hive#2    Inspected by Brain

Brood Chamber;  14 x 12,    QS,  No QC,   4.5  f BAS,  4 f stores, 2+ room,  1 super, Empty.  vc 2, taken out.  Temper 10 Frame with brood on found outside brood nest ,next to stores so moved in. Replaced part drawn frame with shallow for drone trapping.

Hive#3      Inspected by David Parkinson

Brood Chamber:    National Deep.   QNS, No QC,  only 2 f BAS, 1 f stores, room ok,  no supers. temper 10.                       Weak colony.  not building up much, Wax Moth found , may be best to Kill Q and Combine.

Hive#4   Inspected by  David

Brood Chamber:  National Deep and half (shallow).  QNS,  no QC,  8d + 3s f BAS, 3f stores, room ok, 2 supers temp 9   colony building nicely, Healthy hive full of bees.  vb 0. taken out.

Hive#5   Inspected by Anne Chantell Ballard.

Brood Chamber:    National Deep. Shallow added.    QS,  1 QC taken down, not sealed. 7 f BAS,  11 f stores, room ok.     1 super, Temper 8 .   –  cramped in deep so super added to brood chamber.  one deep frame replaced with shallow for drone trapping.   entrance block removed,  Consider an artificial swarm next week if more QC’s

Hive#6   Inspected by Anne

Brood Chamber:   Deep,   Shallow added.  QS,  no QC.  6 f BAS, 3 fd stores,  1 super empty, Temper 9. live wax moth larva on vb,  vb removed  and entrance block removed.

Hive#7   Inspected By Liz Hawes

Brood Chamber:   4 Shallows,  QNS,   10 Qc’s. 2 left on.   20+ f BAS. stores ok. Q excluder  added  under new super of drawn comb,  some full frames of stores moved up from brood chamber into super. replaced by  f drawn comb. will have to do an artificial swarm ASAP.

Hive#8     Inspected By Liz and Mel

Brood Chamber: National Deep.  QS, No QC,  9 f BAS, 2 f stores, room ok, 2 supers. temper 9.  Checked drone brood for varroa, none found, entrance block and vb removed.

Hive#9    Inspected by Richard Williams

Warre Hive.       QS, No Qc’s. 2 f BAS, 1 f stores, room ok,  temper 10.   –  Taped back up several loose frames, very dark and bit mouldy but being used..   better new comb in bottom box.  solid floor cleaned of debris and entrance cleared of old foam.   Small colony.

Hive#10   Inspected by Peter Reader

Top Bar Hive.    QNS, no Qc,   2 sides of BAS,  2 stores,  room ok. temper 10. –  queen has sprung back into action with now 2 sides of brood, but still small colony.   making up slowly.

Nuc  15 F    Not gone through as brood in observation hive at Milland Show.

Nuc 15 H    Inspected by Richard

Brood Chamber: Deep.   QNS, No Qc’s. 3.5 f BAS.  1 f stores, room ok.  Temper 10. rapid feeder taken off,   Moved a further 3 ft, onto concrete platform.

 

Next Apiary Meeting: Sunday 15th May. Lots of things to do

So far the weather forecast is looking promising with sunshine and temperatures of 14 degrees C,  just about workable.     We have to go through the Warre hive as a priority, as it’s not been through yet this year, so should be fun.   Some of you will be going to the Fernhurst revels, so with reduced older members it would be a help,  if some of the other experienced members  who don’t normally attend come to help out .     See you there.

Richard, Liz, and Mel.    Apiary Custodians.

Apiary Report: Sunday 8th May 2016. What a lovely day..

Hot and Sunny, what could go wrong!!!    How about lack of new members, where were you all?   Those that did attend very much enjoyed themselves, with lots of hands on experience opening hives and checking bees.    Things are hotting up in the colonies with most hives showing lots of stores coming in and increasing build up of brood.

Apiary Sundays are a very sociable afternoon, when we can exchange information on bee topics , problems with ones own hives, how to reduce varroa, what to look for in colonies affected with nosema, different hive set ups and general upkeep of hives.

Thirteen members attended Sunday, with only 6 of this years “New Bees”  .  Temperature  in the 20’s C  Sunny.

Hive#1   Inspected by Brian Herbert

Brood Box;    14 x 12     QNS, No QC’s , 4 f brood,BAS. 3 f stores, 2 f room,   1 super,  varroa count 5,  temper 8.              Empty super replaced with another as the frames were not sitting down below the box sides.  Several drone cells were lifted to check for varroa as a few bees were seen with DWV, ( Deformed Wing Virus) thought to be transmitted by high counts of varroa. None were seen on the samples taken.

Hive#2   Inspected by Brian,

Brood Box:    14 x 12 ,   QS,    No QC’s, 3.5 f Brood BAS,  2+ room,  Varroa count 30,  Temper 10.    Dusted with icing sugar for high varroa count.  Overcoming chalk brood problems last week but some diarrhoea noticed on super frames.

Hive#3   Inspected by  Peter Reader.

Brood Box:  National Deep.   QNS,   no QC.   2 f BAS  1.5 f stores,  no supers.  varroa count, Zero. Temper 10                      Still very small colony but building slowly.  good pattern of brood and stores ok. rapid feeder removed.

Hive#4   Inspected by Peter Reader.

Brood Box:    National Deep and Added Shallow.  QS,  no QC,  7 f BAS,  3f stores,   1 super.   no varroa, Temper 9.           Not much room left in deep, so added shallow of DC ( Drawn Comb), big strong colony.

Hive#5    Inspected  by Richard Williams

Brood Box:  National Deep,  QS,  no QC,  5.5 f BAS, 4 f Stores, room ok,  varroa count  3.  Temper 10.                                 Queen located and marked blue as last years. nice and fat still.  lots of stores coming in but not capped yet.

Hive#6   Inspected  by Richard Williams

Brood Box:  National Deep.    QS,  no QC,  5 f BAS, 3 f Stores, room ok,   1 super started filling, no varroa, temper 10.     Some chalk brood on vb,  no mouse droppings as previous weeks, small colony but building up ok.

Hive#7   Inspected by Liz Hawes

Brood Box:  3 All Shallows.( Peter Street)  QNS,  2 QC taken down as damaged, 20 f Brood,  35 f stores, still room ok.   high varroa count at 45.   Temper 9.     Open Queen cells damaged with removal of shallow box as on bottom edge. Lots of bees, strong colony, but bit cramped in shallows.

Hive#8   Inspected by Liz Hawes

Brood Box:  National Deep.   QS,  no QC,  7 f BAS,  stores ok, room ok,  2 supers.  varroa count 2.  Temper 9.                    brood that was in super has now all gone and being filled with stores.

Hive#9     Warre Hive –  Not gone through –  Priority for next week.

Hive#10   Inspected by Peter Reader  Top Bar Hive  Expert..

QS,  no QC, 1 f BAS, stores ok, room ok,    v count , 1.  temper 10.   Moved colony to be central, not making up very well.  same as last week.  50:50 sugar syrup in rapid feeder not taken down, now mouldy so taken off.

Hive#15F    Inspected by Richard

Brood chamber : National deep,  QNS,  no QC.  3.5 f BAS,  1 f stores,  room ok,  no varroa, temper 10.   Rapid feeder taken off ,  making up slowly.

Hive#15H   Inspected by Richard

Poly Nuc, now transfered into National Deep.   QNS, No QC,  3.5 f BAS,  2 f stores,  VC = 2 .  Making up nicely. will be used for display purposes soon,  Moved  3 feet towards concrete slab

 

 

Apiary Next Meeting: Sunday 8th May 14:00 Looking Good for full inspection

Apiary Record: 1st May 2016

IMGP5026

 

Sunday turned out to be lovely and sunny but the wind picked up a bit and it was quite cold at 12 degrees C.   We had 16 members attend, so we managed to go through the hives quite quickly.     “New bees” were paired up with an experienced   member and guided through what to look for in hive inspections,  how to lift and view frames of bees and brood,  checking for queen-right,  moving empty frames into the brood nest to give extra space,  stores and any signs of disease.   All went well but the bees were a bit frisky and a couple of people got there first sting of the season.

All the colonies have so far  come through the winter ok, but No3 , the Warre and the Top Bar hive are all small colonies.    Several were also found to have very little stores, possibly due to the cold spring delaying the flow of nectar.    Feeding with 50 : 50 liquid syrup in rapid feeders were put on as a precaution.

Temperature  12 C,   Sunny.  Breezy. from SW.   Varroa count over 7 days.

 

 

Hive#1    Inspected by Brian Herbert

Brood Chamber:     14 x 12,   Queen Not Seen ( QNS )   4 frames of brood, low stores. lots of drone brood, some chalkbrood  present.  temper 8.   Varroa – zero.    Rapid feeder put on with 50:50 syrup.

Hive#2   Inspected by Brian Herbert

Brood Chamber:    14 x 12,   Queen seen,   3.5 frames of Brood All Stages (BAS)  3 stores, 2 room,  1 super , Temper 6.  Varroa  20.  Some chalk brood present, also some pepper potting brood.

Hive#3   Inspected by David Parkinson.

Brood Chamber:  National  Deep,     QNS,    1.5 f brood  BAS.  only half frame stores. no super, dummy boards either side of brood nest.  No varroa,  Temper 10.  –  very few bees.  Rapid feeder put on with 50:50 syrup.

Hive#4    Inspected by David Parkinson.

Brood Chamber:  National,   QS,    6 frames BAS.  2 stores,  room ok.  1 super.  Varroa- zero.  Temper 9.    Wax moth grub killed. and mouldy brood frame removed.

Hive#5   Inspected by Mel

Brood Chamber:  National,     QNS,   5 f brood,  BAS,  3 f stores,   room ok,  1 super.   Varroa – 1.   Temper  8.                   A  few chalkbrood present.  rapid feeder taken off  as stores ok. some mould in super frame.

Hive#6   Inspected by Mel

Brood Chamber:  National.   QS,   4 f brood BAS ,  1 f stores,  room ok,   No Supers,   Varroa – 2.  Temper 10.                    Few chalkbrood found,  also small colony but building up ok.

Hive#7    Inspected by Liz

Brood Chamber:    National Supers. ( Peter Street Method)  QNS,   13  super frames of Brood  BAS.   13 f stores,     Varroa – 25.   Temper 8.    Chalkbrood on vb,   another super added of drawn comb, as lacking room.

Hive#8    Inspected by Brad Davis

Brood Chamber: National.   QS,   6 f BAS,  2 f stores, 4 f room,  2 supers.  Varroa – 3 .  Some sealed brood still in super above queen excluder,

Hive#9    Warre . Not gone through as too cold for cutting out comb, but very small colony – rapid feeder topped up with 50:50.

Hive#10   Top Bar  Inspected  by  Peter Reader

QS,  1 f brood, BAS,  some stores, room ok,  varroa – 1,  Temper 10.   – Moved colony to be central,  tidy brood pattern. rapid feeder left on.

Hive#15F   Inspected by Brad

Brood Chamber: National,    QS,   3f BAS,   No stores.  plenty room.  Varroa – zero. Temper 10.   Rapid feeder put on with eke.  Nuc colony raised from Queen cell last year.

Hive#15H   Inspected by Brad

Poly Nuc hive  overwintered from last year  and now needed putting into a Brood chamber.   QS,  3 f BAS, 1 f stores,      now into Deep 7 frames of room. varroa -3   Frames transfered into deep and poly nuc dissembled and left for half an hour for the remaining bees to relocate to deep.    Rapid feeder put on with eke.